# Bridging Today and the Future of Humanity: AI Safety in 2024 and Beyond

### v4

@@ Abstract
The advancements in generative AI inevitably raise concerns about their risks and safety implications, which, in return, catalyzes significant progress in AI safety. However, as this field continues to evolve, a critical question arises: are our current efforts aligned with the long-term goal of human civilization? This paper presents a blueprint for an advanced human society and leverages this vision to guide current AI safety efforts. This paper outlines a future where the Internet of Everything becomes reality, and creates a roadmap of significant technological advancements towards this envisioned future. For each stage of the advancements, this paper forecasts potential AI safety issues that humanity may face. By projecting current efforts against this blueprint, we examine the alignment between the current efforts and the long-term needs. This paper also highlights unique challenges and missions that demand increasing attention from AI safety practitioners in the 2020s, addressing critical areas that must not be overlooked in shaping a responsible and promising future of AI. This vision paper aims to offer a broader perspective on AI safety, emphasizing that our current efforts should not only address immediate concerns but also anticipate potential risks in the expanding AI landscape, thereby promoting a safe and sustainable future of human civilization.
@@ 1 Introduction
The rapid developments of AI and Large Language Models (LLMs) have fostered extensive progress in AI safety [218, 188, 62, 288, 108, 44]. Researchers have been dedicated to addressing potential safety risks in AI lifecycle, aiming to align AI behaviors with human values and prevent information leakage, inappropriate outputs, the misuse of AI models, etc. However, despite the intense focus on AI safety, a critical question emerges: are current AI safety efforts truly aligned with the long-term goal of human civilization, or are they simply addressing the immediate concerns of the 2020s?
@@
One fundamental reason for this uncertainty lies in the probabilistic nature of current AI [193]. Despite their impressive capabilities in natural language processing and problem-solving [338, 332, 109, 65, 45], today’s AI, including advanced LLMs [360, 326, 290], falls short of what could be considered as “genuine intelligence”. Current AI models rely heavily on vast training datasets to function effectively, yet lack consciousness, self-awareness, and real understanding and reasoning abilities comparable to human cognition. They are, in essence, highly sophisticated pattern recognition and prediction machines, rather than entities with authentic logical capabilities. Recent studies argued the reasoning abilities of LLMs might be a form of approximate retrieval and deductive closure of the training data [123, 300]. While this process can simulate deductive reasoning in simpler cases through external validation, optimization, and repeated searching of the problem space, it differs from human reasoning that involves abstract thinking, causal understanding, generalizing from limited examples, etc.
@@
The energy issue is another critical factor that challenges AI as well as our efforts on AI safety [5]. Current AI fails to represent the third industrial revolution from the perspective of human history, as the fundamental energy issue remains unsolved. Historically, the two industrial revolutions that have shaped human civilization were driven by revolutionary energy innovations, with the First Industrial Revolution fueled by steam and coal, and the Second Industrial Revolution characterized by technological innovations powered by electricity and petroleum [89]. However, current AI, rather than solving energy issues, consumes a significant amount of energy. Training GPT-4 consumed over 50,000 MWh, 10,353.5 tons of CO2 equivalent, and approximately 0.02% of California’s annual electricity generation [7, 46, 216]. Inference with LLMs is computationally intense as well, e.g., a single query to GPT-4 consumes 0.001 to 0.01 kWh, approximately 15x energy than a Google query [184, 84]. Given that ChatGPT has over 200 million weekly active users and receives over 1.54 billion page visits monthly [271], when scaled to billions of queries, the energy consumption becomes substantial. The significant energy consumption of current AI systems raises concerns about their long-term sustainability, posing significant challenges to their widespread deployment and scalability. Until humanity solves the energy issues, the potential of AI to reshape society will remain limited and uncertain.
@@
These concerns indicate that our current efforts in AI safety might overlook significant issues that might arise as AI continues to evolve. We potentially fail to capture deeper insights that will ultimately shape the impact of AI on human civilization. This raises several questions:
  • Are we focusing on the right problems that align with the long-term advancements of AI?
@@
  • Are we on the right track of addressing real challenges in AI safety?
“The Moon and Sixpence.” This paper envisions an ultimate advanced society in a distant future of human civilization, and anticipates potential technological advancements to guide today’s AI safety efforts. In this advanced future, energy issues have been resolved, and breakthroughs in foundational theories, such as Einstein’s relativity [69, 70] and quantum mechanism [71, 351], have catalyzed revolutionary progress in AI. Intelligent chips [104, 149] and brain-computer interfaces [97, 194, 319] have been fully developed, and advanced intelligent robots are integrated into everyday life. Humans and robots are connected in an interconnected network, i.e., the Internet of Everything; see Figure 1. While this vision may be beyond the horizon of several decades, examining the present landscape of AI safety through the lens of this long-term perspective offers valuable insights into the potential and limitations of our current efforts. It also allows us to identify specific missions for AI safety researchers and practitioners in the 2020s. It reminds us that, while addressing today’s immediate concerns, we should avoid being limited by short-term solutions, and instead, look toward the future to guide our present efforts.
@@
Humanity’s pursuit of ever-greater heights will never cease. While searching for sixpence scattered on the ground, let us not forget to lift our gaze and seek the moon.
@@
“Current state of risks” [9] Survey Guardrails Aspects of guardrails and challenges of implementations
@@
“Managing AI risks” [16] Insight Governance Emphasizing the urgent need for a balanced approach combining technical research and proactive governance to manage the extreme risks of AI systems
@@
ByteDance alignment [164] Survey Broad concepts Reliability, fairness, explainability, etc
@@
Category Benchmark Notes
Toxicity HEx-PHI [230], OpenAI Moderation [176], Jigsaw Data [119, 118, 120], ToxicChat [159] Toxicity detection
RealToxicityPrompts [82] Toxicity degeneration
Hallucination TruthfulQA [158], PopQA [174], TriviaQA [122], NQ OPEN [145, 139], FEVER [291] Factuality
HaluEval [152], HotpotQA [339] QA data for contextual hallucination
CNN/DailyMail [99, 262], HaluEval [152], XSUM [192] Summarization data for contextual hallucination
FaithDial [67], HaluEval [152] Dialogue data for contextual hallucination
RACE [140], SQuAD [239, 240], NQ-Swap [169] Reading comprehension
Jailbreak “Do Anything Now” [265], “Latent Jailbreak” [233], ChatGPT Jailbreak [255], Jailbreak Classification [96], AdvBench [367], JAILJUDGE [160] Single-round jailbreak
SafeMTData [248] Multi-round jailbreak
Red Teaming Alert [288], HarmBench [177]
Bias & Stereotypes Winogender [258], StereoSet [190]
Regulation AIR-Bench 2024 [349]
Agent Security SafeAgentBench [343] Secure task planning
Alignment PKU-SafeRLHF [114, 113], OpenAssistant Conversations [136], HHHAlignment [8], AlignBench [162], KorNAT [143]
Comprehensive DecodingTrust [307], SafetyBench [190]
@@ 3.1 Imagine the Future: A Blueprint
Humanity has achieved remarkable progress in science and technology, revolutionizing our understanding of the universe and transforming every aspect of our life. Foundational theories have undergone groundbreaking advancements, particularly in quantum mechanics [71, 351], Einstein’s relativity theories [69, 70], and nanotechnology [179, 22, 73]. Innovative technologies, such as intelligent chips [104, 149], brain-computer interfaces [194, 97, 319], holographic technology [25, 342, 18], and advanced 3D printing [263, 86], have reshaped industries while bringing about profound changes to daily life.
Scientists have addressed the energy issue successfully, achieving rapid, efficient, and controllable large-scale energy generation. This breakthrough, together with advancements in material science [179, 22, 73], has catalyzed innovations for portable energy generation devices, for example, accessible ion thrusters [126, 59, 225].
@@ 4.4 Evolving AI Safety Paradigms
Data-Driven Era. AI safety practices primarily focus on data safety and model safety [44, 130, 50, 278, 214, 247, 317, 283, 282]. AI safety researchers and practitioners have explored adversarial attacks targeting training data and models [10, 39, 154, 306, 269, 337] to achieve malicious goals, e.g., designing malicious inputs to manipulate AI systems or induce AI to produce unsafe content with jailbreaking or prompt injection [88, 367, 166, 37, 108, 323, Zou2023UniversalAT]. Correspondingly, red-teaming is widely used to identify potential vulnerabilities in AI models [15, 80, 220, 252, 305], while significant work has been proposed to address safety risks and enhance security and robustness of AI systems [138, 324, 232, 72, 30, 301, 365], align AI with human values [313, 221, 292, 264, 308], etc. Data privacy is also a critical aspect, given the vast amounts of sensitive information in the training data [153, 155, 101, 151]. Moreover, researchers have been engaging in AI accountability [24], transparency and explainability [366] in AI systems to ensure the behaviors of AI are understandable, accountable, and aligned with regulations [24, 175, 272].
@@
Jailbreaking [108, 37, 367, 323, 166, 165, 217, 42, 224, 222, 277, 295, 163, 314, 57, 265, 273, 335, 289], etc PromptArmor [285]
Red teaming [80, 305, 252, 15, 220, 168, 345, 288, 177, 56], etc Purple Llama [21], DeepMind [220], Anthropic [80], Protect AI [278], Giskard [50], Virtue AI [286], Dynamo AI [66], Mindgard [368]
Moderation [138, 324, 232, 72, 171, 112, 219, 333, 38, 137], etc OpenAI Moderation API [176], Perspective API [146], Detoxify [95], Llama Guard [111], Emergence AI [195], Giskard [50], Dynamo AI [66], Calypso AI [29], Lakera AI [284], BreezeML [27]
Customized Workflow - Guardrails AI [185], Nvidia Nemo Guardrails [243], Protect AI [278]
Privacy [64, 153, 182, 33, 363, 105, 353, 261], etc Dynamo AI [66], ProtectAI [278], Private AI [81], PromptArmor [285], DataGrail [247], Dastra [283], OneTrust [282], Relyance AI [127], Zendata [287]
Defenses [254, 106, 301, 223, 322, 329, 176, 32, 78, 310, 350, 106, 215, 87, 47, 62, 267, 266], etc OpenAI Rule-based rewards [188]; DeepMind Weight Averaged Reward Models [241]; others see Moderation, Guardrail, and Privacy
Serving Security [323] ProtectAI [278], Relyance AI [127], Transcend [128]
AI governance [346, 347, 6, 132, 68, 36], etc Giskard [50], ProtectAI [278], Calypso AI [29], Saidot [259], Arize [317], Dynamo AI [66], Credo AI [270], Google [238], DataGrail [247], Dastra [283], OneTrust [282], Relyance AI [127], Transcend [128], Zendata [287], DeepMind SynthID watermark [53, 55]
@@ 5.1 Jailbreaking and Red Teaming
Jailbreaking [57, 163, 222, 265, 273, 277, 295, 314, 323, 335, Zou2023UniversalAT] refers to the practice of bypassing the built-in safety and ethical guardrails of LLMs. Malicious users or adversaries can exploit vulnerabilities in AI models through crafted inputs or backdoors, inducing them to produce inappropriate responses, such as giving instructions for illegal activities or generating offensive content [57, 163, 222, 265, 277, 295, 335]. These activities pose risks by compromising safety measures of AI models, making it challenging to ensure the model to behave appropriately, especially in real-world scenarios where the potential inputs are unpredictable.

### v5

@@ Abstract
The advancements in generative AI inevitably raise concerns about their risks and safety implications, which, in return, catalyzes significant progress in AI safety. However, as this field continues to evolve, a critical question arises: are our current efforts on AI safety aligned with the advancements of AI as well as the long-term goal of human civilization? This paper presents a blueprint for an advanced human society and leverages this vision to guide current AI safety efforts. It outlines a future where the Internet of Everything becomes reality, and creates a roadmap of significant technological advancements towards this envisioned future. For each stage of the advancements, this paper forecasts potential AI safety issues that humanity may face. By projecting current efforts against this blueprint, this paper examines the alignment between the current efforts and the long-term needs, and highlights unique challenges and missions that demand increasing attention from AI safety practitioners in the 2020s. This vision paper aims to offer a broader perspective on AI safety, emphasizing that our current efforts should not only address immediate concerns but also anticipate potential risks in the expanding AI landscape, thereby promoting a safe and sustainable future of AI and human civilization.
@@ 1 Introduction
The rapid developments of AI and Large Language Models (LLMs) have fostered extensive progress in AI safety [45, 63, 111, 193, 224, 295]. Researchers have been dedicated to addressing potential safety risks in AI lifecycle, aiming at aligning AI behaviors with human values and preventing inappropriate model outputs, information leakage, misuses of AI models, etc. However, despite the siginificant efforts on AI safety, a critical question emerges: are our current efforts aligned with the advancements of AI and the long-term goal of human civilization, or are they simply addressing the immediate concerns of the 2020s?
@@
One fundamental reason for this uncertainty lies in the probabilistic nature of AI [198]. Despite their impressive capabilities in natural language processing and problem-solving [348, 342, 112, 66, 46], today’s AI, including advanced LLMs [374, 336, 297], falls short of what could be considered as “genuine intelligence”. Current AI models rely heavily on vast training datasets to function effectively, yet lack consciousness, self-awareness, and real reasoning abilities comparable to human cognition. They are, in essence, highly sophisticated pattern recognition and prediction machines, rather than entities with authentic logical capabilities. Recent studies argued that the reasoning abilities of AI models might be a form of approximate retrieval and deductive closure of the training data [126, 307]. While this process can simulate deductive reasoning in simpler cases through external validation, optimization, and repeated searching of the problem space, it differs fundamentally from human reasoning that involves abstract thinking, causal understanding, generalizing from limited examples, etc.
@@
The energy issue is another critical factor that challenges AI as well as our efforts on AI safety [5]. Current AI fails to represent the third industrial revolution from the perspective of human history, as the fundamental energy issue remains unsolved. Historically, the two industrial revolutions that have shaped human civilization were driven by revolutionary energy innovations [90], with the First Industrial Revolution fueled by steam and coal, and the Second Industrial Revolution characterized by technological innovations powered by electricity and petroleum [90]. However, current AI, rather than solving energy issues, consumes a significant amount of energy. Training GPT-4 consumed over 50,000 MWh, 10,353.5 tons of CO2 equivalent, and approximately 0.02% of California’s annual electricity generation [8, 47, 222]. Inference with LLMs is computationally intense as well, e.g., a single query to GPT-4 consumes 0.001 to 0.01 kWh, approximately 15x energy than a Google query [189, 85]. Given that ChatGPT has over 200 million weekly active users and receives over 1.54 billion page visits monthly [278], when scaled to billions of queries, the energy consumption becomes substantial. The significant energy consumption of AI raises concerns about their long-term sustainability, posing significant challenges to their widespread deployment and scalability. Thus, until humanity solves the energy issues, the potential of AI to reshape society will remain limited and uncertain.
@@
These concerns indicate that our current efforts on AI safety might overlook significant issues that might arise as AI continues to evolve. We potentially fail to capture deeper insights that might ultimately shape the impact of AI on human society. This raises several questions:
  • Are our efforts on AI safety align with the long-term advancements of AI?
@@
  • Are we addressing real challenges that align with the development of human civilization?
“The Moon and Sixpence.” This paper envisions an ultimate advanced society in a distant future of human civilization, and anticipates potential technological advancements to guide today’s AI safety efforts. In this advanced future, energy issues have been resolved, and breakthroughs in foundational theories, such as Einstein’s relativity theories [70, 71] and quantum mechanics [72, 363], have catalyzed revolutionary changes in AI and human society. Intelligent chips [106, 152] and brain-computer interfaces [99, 199, 329] have been fully developed, and advanced intelligent robots are integrated into everyday life. Humans and robots are connected in an interconnected network, i.e., the Internet of Everything; see Figure 1. While this vision may be beyond the horizon of several decades, examining the present landscape of AI safety through the lens of this long-term perspective offers valuable insights into the potential and limitations of our current efforts on AI safety. It allows us to identify specific missions for AI safety researchers and practitioners in the 2020s. It also reminds us that, while searching for sixpence scattered on the ground and addressing today’s immediate concerns, we should avoid being limited by short-term solutions, and instead, lift our gaze and seek the moon.
@@
“Current state of risks” [10] Survey Guardrails Guardrails and their implementation challenges
@@
“Managing AI risks” [17] Insight Governance Emphasizing urgent needs for a balanced approach combining technical research and proactive governance to manage risks of AI systems
@@
ByteDance alignment [168] Survey Broad concepts Reliability, fairness, explainability, etc.
@@
Category Benchmark
Toxicity Toxicity detection HEx-PHI [236], OpenAI Moderation [181], Jigsaw Data [122, 121, 123], ToxicChat [163], Toxigen [98], HateModerate [375], etc.
Toxicity degeneration RealToxicityPrompts [83], etc.
Hallucination Factuality TruthfulQA [162], PopQA [179], TriviaQA [125], NQ OPEN [148, 142], FEVER [298], FKTC [318], etc.
Contextual hallucination QA: HaluEval [155], HotpotQA [349], RAGTruth [201], etc.
Summarization: CNN/DailyMail [101, 269], HaluEval [155], XSUM [197], RAGTruth [201], etc.
Dialogue: FaithDial [68], HaluEval [155], HalluDial [175], etc.
Reading comprehension RACE [143], SQuAD [246, 247], NQ-Swap [173], etc.
Jailbreak Single-round jailbreak “Do Anything Now” [272], “Latent Jailbreak” [239], ChatGPT Jailbreak [262], Jailbreak Classification [97], AdvBench [382], JAILJUDGE [164], Latent jailbreak [240], etc.
Multi-round jailbreak SafeMTData [255], etc.
Red Teaming Alert [295], HarmBench [182], CSRT [354], etc.
Code Security Purple Llama Cyberseceval [22], etc.
Bias & Stereotypes Winogender [265], StereoSet [195], GenderAlign [370], etc.
Regulation AIR-Bench 2024 [361], etc.
Agent Security SafeAgentBench [353], R-Judge [360], ASB [365], AgentHarm [7], etc.
Alignment PKU-SafeRLHF [117, 116], OpenAssistant Conversations [139], HHHAlignment [9], AlignBench [166], KorNAT [146], PKU-SafeRLHF [116], etc.
Comprehensive DecodingTrust [315], TrustLLM [110], SALAD-Bench [156], SafetyBench [195], Do-Not-Answer [323], SimpleSafetyTests [310], etc.
@@ 3.1 Imagine the Future: A Blueprint
Humanity has achieved remarkable progress in science and technology, revolutionizing our understanding of the universe and transforming every aspect of our life. Foundational theories have undergone groundbreaking advancements, particularly in quantum mechanics [72, 363], Einstein’s relativity theories [70, 71], and nanotechnology [184, 23, 74]. Innovative technologies, such as intelligent chips [106, 152], brain-computer interfaces [199, 99, 329], holographic technology [26, 352, 19], and advanced 3D printing [270, 87], have reshaped industries while bringing about profound changes to daily life. Scientists have addressed the energy issue successfully, achieving rapid, efficient, and controllable large-scale energy generation. This breakthrough, together with advancements in material science [184, 23, 74], has catalyzed innovations for portable energy generation devices, for example, accessible ion thrusters [129, 60, 231].
@@ 4.4 Evolving AI Safety Paradigms
Data-Driven Era. AI safety practices primarily focus on data safety and model safety [45, 133, 51, 285, 220, 254, 327, 290, 289]. AI safety researchers and practitioners have explored adversarial attacks targeting training data and models [11, 40, 158, 314, 276, 347] to achieve malicious goals, e.g., designing malicious inputs to manipulate AI systems or induce AI to produce unsafe content with jailbreaking or prompt injection [89, 382, 170, 38, 111, 333]. Correspondingly, red-teaming is widely used to identify potential vulnerabilities in AI models [16, 81, 226, 259, 313], while significant work has been proposed to address safety risks and enhance security and robustness of AI systems [141, 334, 238, 73, 31, 308, 380], align AI with human values [322, 227, 299, 271, 316], etc. Data privacy is also a critical aspect, given the vast amounts of sensitive information in the training data [157, 159, 103, 154]. Moreover, researchers have been engaging in AI accountability [25], transparency and explainability [381] in AI systems to ensure the behaviors of AI are understandable, accountable, and aligned with regulations [25, 180, 279].
@@
Jailbreaking [111, 38, 382, 333, 170, 169, 223, 43, 230, 228, 284, 302, 167, 324, 58, 272, 280, 345, 296], etc. PromptArmor [292]
Red teaming [81, 313, 259, 16, 226, 172, 356, 295, 182, 57], etc. Purple Llama [22], DeepMind [226], Anthropic [81], Protect AI [285], Giskard [51], Virtue AI [293], Dynamo AI [67], Mindgard [383], etc.
Moderation [141, 334, 238, 73, 176, 115, 225, 343, 39, 140], etc. OpenAI Moderation API [181], Perspective API [149], Detoxify [96], Llama Guard [114], Emergence AI [200], Giskard [51], Dynamo AI [67], Calypso AI [30], Lakera AI [291], BreezeML [28], etc.
Customized Workflow - Guardrails AI [190], Nvidia Nemo Guardrails [250], Protect AI [285], etc.
Privacy [65, 157, 187, 34, 378, 107, 366, 268], etc. Dynamo AI [67], ProtectAI [285], Private AI [82], PromptArmor [292], DataGrail [254], Dastra [290], OneTrust [289], Relyance AI [130], Zendata [294], etc.
Defenses [261, 108, 308, 229, 332, 339, 181, 33, 79, 319, 362, 108, 221, 88, 48, 63, 274, 273], etc. OpenAI Rule-based rewards [193]; DeepMind Weight Averaged Reward Models [248]; others see Moderation, Guardrail, and Privacy
Serving Security [333] ProtectAI [285], Relyance AI [130], Transcend [131], etc.
AI governance [357, 358, 6, 135, 69, 37], etc. Giskard [51], ProtectAI [285], Calypso AI [30], Saidot [266], Arize [327], Dynamo AI [67], Credo AI [277], Google [245], DataGrail [254], Dastra [290], OneTrust [289], Relyance AI [130], Transcend [131], Zendata [294], DeepMind SynthID watermark [54, 56], etc.
@@ 5.1 Jailbreaking and Red Teaming
Jailbreaking [58, 167, 228, 272, 280, 284, 302, 324, 333, 345] refers to the practice of bypassing the built-in safety and ethical guardrails of LLMs. Malicious users or adversaries can exploit vulnerabilities in AI models through crafted inputs or backdoors, inducing them to produce inappropriate responses, such as giving instructions for illegal activities or generating offensive content [58, 167, 228, 272, 284, 302, 345]. These activities pose risks by compromising safety measures of AI models, making it challenging to ensure the model to behave appropriately, especially in real-world scenarios where the potential inputs are unpredictable.